CH: CH, OH CH2=CCH3 + Cl2 Part D CH,OH CH,CH=CHCH, + Bry Part C CH,CH=CHCH, + HCI-> Part B CH=CHCH3 + 2NaI + HBr +

for each of these it says to draw the major product for each reaction. a.b.c.d parts. pls pls help
Answer #1

Иy (H3 CH₂ = C CH2 + Cl₂ – CH = C- CH₂ choch & CH ₂ CH = CH CHz ABY - CH₂ CH - CH CH3 3х осна ен,си си, — си си- скасы -cң, c